# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

### v1

@@ Appendix A Summary of Notation
$T$; $t\in[T]$ Total queries (horizon $T$); $t$ is query index in $[T]=\{1,\dots,T\}$.
@@
$\tau_{0}$ Initial time for the algorithm, typically 00.
$\mathcal{A}$ The online LLM routing algorithm/policy.
@@
$\alpha_{m}$ Capacity limit (maximum load share / selection probability constraint) for LLM $m$.
@@
Appendix B Technical Analysis
@@ B.1 Concentration Inequality
Lemma 2 ([9], adapted).
Let $\mathcal{D}_{k}(\mathcal{A})$ be the set of deployed models in stage $k$. For any vectors $\boldsymbol{a}=(a_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ and $\boldsymbol{n}=(n_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ where $a_{m},n_{m}\geq 0$,
@@
Follows from applying Lemma 1 to the sequence of observed performances $r_{s}$ (for $m_{s}=m$) and observed costs $c_{s}$ (for $m_{s}=m$). For a fixed model $m$, the rewards $r_{s}$ (when $m_{s}=m$) are i.i.d. samples from $\nu_{m}(\cdot)$ with mean $\mu_{m}$. Similarly, costs $c_{s}$ (when $m_{s}=m$) are effectively i.i.d. samples with mean $\mathbb{E}[c_{m}]$. Thus, the conditional expectation $\mathbb{E}[x_{i}|x_{1},\ldots,x_{i-1}]$ in Lemma 1 becomes the true mean $\mu_{m}$ (or $\mathbb{E}[c_{m}]$). The derivation is analogous to Lemma 4.3 of [9]. For instance, for performance:
@@
B.2 Regret Decomposition with Time-Varying Benchmark
@@
For a given stage $k$ (time interval $[\tau_{k},\tau_{k+1})$ of length $T_{k}=\tau_{k+1}-\tau_{k}$) where StageRoute (during its Model Deployment Phase) deploys the set $\mathcal{D}_{k}=\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$, $V(b,\mathcal{D}_{k})$ represents the optimal expected reward per query achievable using only models from the deployed set $\mathcal{D}_{k}$. The total optimal expected performance within this stage using $\mathcal{D}_{k}$ is $\text{OPT}_{k}=T_{k}\cdot V(b,\mathcal{D}_{k})$. Note that $\mathcal{D}_{k}$ is determined by StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates available at $\tau_{k}$. Thus, $\mathcal{D}_{k}$ and consequently $V(b,\mathcal{D}_{k})$ (and $\text{OPT}_{k}$) are random variables, dependent on the algorithm’s choices and observations up to time $\tau_{k}$.
@@
B.3 Analysis of Deployment Regret
@@
    Parameter Uncertainty: Inaccurate estimates of model performance ($\mu_{m}$) and cost ($\mathbb{E}[c_{m}]$) may result in suboptimal deployment decisions. This source corresponds to models that have already been deployed in one or more of the previous $k-1$ stages.
@@
    Model Discovery Bottleneck: The constraint that at most $M_{\max}$ models can be deployed concurrently may exclude promising but underexplored models—particularly newly arrived ones—from being included in $\mathcal{D}_{k}$. This prevents timely evaluation and utilization, contributing to additional regret. This case pertains to models that have not been selected in any of the preceding $k-1$ stages.
@@
For the analysis, we recall that $V(b,\mathcal{S})$ is the optimal performance rate for a model set $\mathcal{S}$ (defined in Section 3). Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate achievable using all models available at the start of stage $k$, and $V_{k}=V(b,\mathcal{D}_{k})$ be the optimal rate achievable using the subset $\mathcal{D}_{k}\subseteq\mathcal{M}_{\tau_{k}}$ selected by StageRoute. $V^{*}_{k}$ is deterministic given $\tau_{k}$, while $V_{k}$ (through $\mathcal{D}_{k}$) is a random variable. A key assumption for bounding the deployment regret due to parameter uncertainty involves the Lagrange multipliers associated with the budget constraint in the definition of $V^{*}_{k}$. Let $\lambda^{*}_{k}\geq 0$ be this optimal Lagrange multiplier. We assume that $\lambda^{*}_{k}$ is uniformly bounded by a constant $\Lambda$ for all $k$. This is a common assumption in the analysis of learning algorithms with budget constraints and is often justified when standard regularity conditions (such as Slater’s condition, which we assume in Assumption 1) hold for the underlying optimization problems, particularly given that our problem parameters (rewards, costs, capacities $\alpha_{m}$) are bounded and expected costs $\mathbb{E}[c_{m}]$ are lower-bounded by $c_{1}>0$.
@@
Let $\mathcal{E}=\cap_{k=1}^{K}\mathcal{E}_{k}$. By a union bound over all $N$ models in the universe $\mathcal{M}_{T}$ and $K$ stages, $\mathbb{P}(\mathcal{E})\geq 1-\delta$. We condition the analysis on $\mathcal{E}$. (Note: $\mu^{U}_{m}$ is $\bar{\mu}_{m}(\tau_{k})+2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ and $c^{L}_{m}$ is $\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ as per Eq. (6,7) in the algorithm description, projected onto $[0,1]$ and $[c_{1},c_{2}]$ respectively. The inequalities above capture the desired properties on the good event $\mathcal{E}_{k}$).
@@
Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate with available models at stage $k$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$ be the set selected by the Model Deployment Phase of StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates at $\tau_{k}$, via solution $d^{*}$ (from Eq. (3)). Let $V_{k}=V(b,\mathcal{D}_{k})$. On the good event $\mathcal{E}_{k}$, the deployment gap for stage $k$ is bounded as:
@@
Let $d^{\textnormal{opt},k}$ be an optimal solution achieving $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$. Let $d^{*}$ be the solution found by StageRoute’s Model Deployment Phase (using Eq. (3)) at $\tau_{k}$ when optimizing over $\mathcal{M}_{\tau_{k}}$ using $\mu^{U}_{m}(\tau_{k})$ and $c^{L}_{m}(\tau_{k})$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$. Let $V_{k}=V(b,\mathcal{D}_{k})$.
On the event $\mathcal{E}_{k}$, the confidence bounds hold for all $m\in\mathcal{M}_{\tau_{k}}$. Specifically, $\mu_{m}\leq\mu^{U}_{m}(\tau_{k})\leq\mu_{m}+\text{rad}_{\mu}(m,\tau_{k})$ and $\mathbb{E}[c_{m}]-\text{rad}_{c}(m,\tau_{k})\leq c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$. The true optimal solution $d^{\textnormal{opt},k}$ for the set $\mathcal{M}_{\tau_{k}}$ satisfies $\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$, uses $\leq M_{\max}$ models from $\mathcal{M}_{\tau_{k}}$, etc. Since $c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$ on $\mathcal{E}_{k}$, we have $\sum_{m\in\mathcal{M}_{\tau_{k}}}c^{L}_{m}(\tau_{k})d^{\textnormal{opt},k}_{m}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$. Thus, $d^{\textnormal{opt},k}$ is a feasible solution for the optimization problem solved by StageRoute (Eq. (3) applied to $\mathcal{M}_{\tau_{k}}$).
@@
Substituting these into Eq. (11):
@@
This bounds the true performance of the distribution $d^{*}$ chosen by the algorithm. Now we relate this to $V_{k}=V(b,\mathcal{D}_{k})$, the optimal performance within the chosen set $\mathcal{D}_{k}=\textnormal{supp}(d^{*})\subseteq\mathcal{M}_{\tau_{k}}$. The distribution $d^{*}$ is supported on $\mathcal{D}_{k}$ and uses at most $M_{\max}$ models (due to the constraint in Eq. (3)). We examine its feasibility w.r.t. the true budget constraint. On $\mathcal{E}_{k}$:
@@
The term $d^{*}_{k,m}$ represents the selection weight for model $m$ in the deployment optimization at stage $k$. Thus, $\sum_{m\in\mathcal{D}_{k}}C_{m}(\tau_{k})d^{*}_{k,m}$ is a weighted average of the combined confidence radii for the deployed models. We are considering the case where the model has already been selected in the previous $k-1$ stages. Thus, we can bound the sum of $T_{k}(V^{*}_{k}-V_{k})$ by terms related to $\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}C_{m}(\tau_{k})$. Specifically,
@@
The StageRoute algorithm employs UCBs for rewards ($\mu^{U}_{m}$) and LCBs for costs ($c^{L}_{m}$) in its DeployOPT phase (Eq. (3)). For a model $m$ that is newly available at stage $k$ (i.e., $t_{m}\leq\tau_{k}$ and its count of previous selections $N_{m}(\tau_{k})=0$), its initial empirical averages $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ are set based on priors. The confidence radius $f_{rad}(v,N_{m}(\tau_{k})+1)=\sqrt{\frac{\gamma v}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}$ becomes large for $N_{m}(\tau_{k})=0$. Specifically, with $N_{m}(\tau_{k})+1=1$:
@@
The total deployment regret bound in Lemma 8 highlights two distinct challenges in the deployment phase. The terms $\mathcal{O}(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})})$ and $\mathcal{O}(M_{\max}K\log(NT/\delta))$ capture the cost of learning the parameters of models that are considered for deployment. This cost depends on the horizon $T$, the number of deployment slots $M_{\max}$, the number of stages $K$, and logarithmic factors related to the total number of models $N$ and confidence $\delta$. The term $\mathcal{O}(NT/(M_{\max}K))$ reflects the structural cost imposed by the discovery bottleneck: when the universe of models $N$ is large compared to the deployment capacity $M_{\max}$ and the number of adaptation opportunities $K$, there is an inherent regret incurred in sequentially exploring models to identify the best ones. This term can dominate if $N$ is very large or $K$ is small, underscoring the importance of the deployment frequency and capacity in dynamic LLM environments.
B.4 Bounding the Total Routing Regret
We now bound the total routing regret term $\mathcal{R}_{\textnormal{routing}}(T)$ identified in Lemma 4. This involves summing the per-stage regrets incurred by the Request Routing Phase of StageRoute due to using estimated model parameters within the deployed sets $\mathcal{D}_{k}$. This part of the proof follows a similar structure to the analysis of UCB-style algorithms for the Bandits with Knapsacks problem [5].
@@ B.4 Bounding the Total Routing Regret
The proof follows the structure of Lemma 4.4 in [9], adapted to our notation for models $m\in\mathcal{D}_{k}(\mathcal{A})$ and performance $\mu_{m}$, UCB $\mu^{U}_{m}(t)$, chosen model $m_{t}$, routing distribution $p_{t}^{*}$ (from Eq. (7)), and summing over $t\in[\tau_{k},\tau_{k+1}-1]$ (length $T_{k}$).
@@
$\leq\mathcal{O}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\sum_{\text{plays }j\text{ of }m\text{ in stage }k}f_{rad}(\hat{\mu}_{m}(\text{at play }j),N_{m}(\text{at play }j)+1)\right)$
@@
The proof mirrors that of Lemma 9 (and Lemma 4.5 in [9]), replacing performance with cost, $\mu_{m}$ with $\mathbb{E}[c_{m}]$, $\mu^{U}_{m}(t)$ with $c^{L}_{m}(t)$, and $r_{t}$ with $c_{t}$. The probability distribution $p_{t}^{*}$ is from Eq. (7). Key steps involve bounding $|\sum c_{t}-\sum\mathbb{E}[c_{m_{t}}]|$, $|\sum(\sum_{m}c^{L}_{m}(t)p_{t}^{*}(m))-\sum c^{L}_{m_{t}}(t)|$, and $|\sum\mathbb{E}[c_{m_{t}}]-\sum c^{L}_{m_{t}}(t)|$. The last term is bounded similarly using Lemma 3 and Lemma 2:
@@
From the algorithm’s choice of $p_{t}^{*}$ (solving Eq. (7)) and property (1):
@@
Taking expectation conditioned on $\mathcal{D}_{k}$ (and implicitly on $\mathcal{E}$ for the bounds to hold), the result follows. The assumption $M_{k}\gamma\leq\mathcal{O}(B_{k})$ is used in concentration bounds for costs. ∎
@@
B.5 Total Regret Bound

### v2

@@ Appendix B Summary of Notation
$T$; $t\in[T]$ Total queries (horizon $T$); $t$ is query index in $T]=\{1,\dots,T\}$.
@@
$\tau_{0}$ Initial time for the algorithm, typically 0.
$\mathcal{A}$ The online deployment and routing algorithm.
@@
$\alpha_{m}$ Throughput limit (maximum load share / selection probability constraint) for LLM $m$.
@@
Appendix C Technical Analysis
@@ C.1 Concentration Inequality
Lemma 2 (Babaioff et al. (2015), adapted).
Let $\mathcal{D}_{k}(\mathcal{A})$ be the set of deployed models in stage $k$. For any vectors $\bm{a}=(a_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ and $\bm{n}=(n_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ where $a_{m},n_{m}\geq 0$,
@@
Follows from applying Lemma 1 to the sequence of observed performances $r_{s}$ (for $m_{s}=m$) and observed costs $c_{s}$ (for $m_{s}=m$). For a fixed model $m$, the rewards $r_{s}$ (when $m_{s}=m$) are i.i.d. samples from $\nu_{m}(\cdot)$ with mean $\mu_{m}$. Similarly, costs $c_{s}$ (when $m_{s}=m$) are effectively i.i.d. samples with mean $\mathbb{E}[c_{m}]$. Thus, the conditional expectation $\mathbb{E}[x_{i}|x_{1},\ldots,x_{i-1}]$ in Lemma 1 becomes the true mean $\mu_{m}$ (or $\mathbb{E}[c_{m}]$). The derivation is analogous to Lemma 4.3 of Babaioff et al. (2015). For instance, for performance:
@@
C.2 Regret Decomposition with Time-Varying Benchmark
@@
For a given stage $k$ (time interval $\tau_{k},\tau_{k+1})$ of length $T_{k}=\tau_{k+1}-\tau_{k}$) where StageRoute (during its Model Deployment Phase) deploys the set $\mathcal{D}_{k}=\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$, $V(b,\mathcal{D}_{k})$ represents the optimal expected reward per query achievable using only models from the deployed set $\mathcal{D}_{k}$. The total optimal expected performance within this stage using $\mathcal{D}_{k}$ is $\text{OPT}_{k}=T_{k}\cdot V(b,\mathcal{D}_{k})$. Note that $\mathcal{D}_{k}$ is determined by StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates available at $\tau_{k}$. Thus, $\mathcal{D}_{k}$ and consequently $V(b,\mathcal{D}_{k})$ (and $\text{OPT}_{k}$) are random variables, dependent on the algorithm’s choices and observations up to time $\tau_{k}$.
@@
C.3 Analysis of Deployment Regret
@@
    Parameter Uncertainty: Inaccurate estimates of model performance ($\mu_{m}$) and cost ($\mathbb{E}[c_{m}]$) may result in suboptimal deployment decisions. This source corresponds to models that have already been deployed in one or more of the previous k1k-1 stages.
@@
    Model Discovery Bottleneck: The constraint that at most $M_{\max}$ models can be deployed concurrently may exclude promising but underexplored models—particularly newly arrived ones—from being included in $\mathcal{D}_{k}$. This prevents timely evaluation and utilization, contributing to additional regret. This case pertains to models that have not been selected in any of the preceding k1k-1 stages.
@@
For the analysis, we recall that $V(b,\mathcal{S})$ is the optimal performance rate for a model set $\mathcal{S}$ (defined in Section 2). Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate achievable using all models available at the start of stage $k$, and $V_{k}=V(b,\mathcal{D}_{k})$ be the optimal rate achievable using the subset $\mathcal{D}_{k}\subseteq\mathcal{M}_{\tau_{k}}$ selected by StageRoute. $V^{*}_{k}$ is deterministic given $\tau_{k}$, while $V_{k}$ (through $\mathcal{D}_{k}$) is a random variable. A key assumption for bounding the deployment regret due to parameter uncertainty involves the Lagrange multipliers associated with the budget constraint in the definition of $V^{*}_{k}$. Let $\lambda^{*}_{k}\geq 0$ be this optimal Lagrange multiplier. We assume that $\lambda^{*}_{k}$ is uniformly bounded by a constant $\Lambda$ for all $k$. This is a common assumption in the analysis of learning algorithms with budget constraints and is often justified when standard regularity conditions (such as Slater’s condition, which we assume in Assumption 1) hold for the underlying optimization problems, particularly given that our problem parameters (rewards, costs, $\alpha_{m}$) are bounded and expected costs $\mathbb{E}[c_{m}]$ are lower-bounded by $c_{1}>0$.
@@
Let $\mathcal{E}=\cap_{k=1}^{K}\mathcal{E}_{k}$. By a union bound over all $N$ models in the universe $\mathcal{M}_{T}$ and $K$ stages, $\mathbb{P}(\mathcal{E})\geq 1-\delta$. We condition the analysis on $\mathcal{E}$. (Note: $\mu^{U}_{m}$ is $\bar{\mu}_{m}(\tau_{k})+2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ and $c^{L}_{m}$ is $\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ as per Eq. (6,7) in the algorithm description, projected onto $[0,1]$ and $c_{1},c_{2}]$ respectively. The inequalities above capture the desired properties on the good event $\mathcal{E}_{k}$).
@@
Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate with available models at stage $k$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$ be the set selected by the Model Deployment Phase of StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates at $\tau_{k}$, via solution $d^{*}$ (from Eq. equation 3). Let $V_{k}=V(b,\mathcal{D}_{k})$. On the good event $\mathcal{E}_{k}$, the deployment gap for stage $k$ is bounded as:
@@
Let $d^{\textnormal{opt},k}$ be an optimal solution achieving $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$. Let $d^{*}$ be the solution found by StageRoute’s Model Deployment Phase (using Eq. equation 3) at $\tau_{k}$ when optimizing over $\mathcal{M}_{\tau_{k}}$ using $\mu^{U}_{m}(\tau_{k})$ and $c^{L}_{m}(\tau_{k})$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$. Let $V_{k}=V(b,\mathcal{D}_{k})$.
On the event $\mathcal{E}_{k}$, the confidence bounds hold for all $m\in\mathcal{M}_{\tau_{k}}$. Specifically, $\mu_{m}\leq\mu^{U}_{m}(\tau_{k})\leq\mu_{m}+\text{rad}_{\mu}(m,\tau_{k})$ and $\mathbb{E}[c_{m}]-\text{rad}_{c}(m,\tau_{k})\leq c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$. The true optimal solution $d^{\textnormal{opt},k}$ for the set $\mathcal{M}_{\tau_{k}}$ satisfies $\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$, uses $\leq M_{\max}$ models from $\mathcal{M}_{\tau_{k}}$, etc. Since $c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$ on $\mathcal{E}_{k}$, we have $\sum_{m\in\mathcal{M}_{\tau_{k}}}c^{L}_{m}(\tau_{k})d^{\textnormal{opt},k}_{m}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$. Thus, $d^{\textnormal{opt},k}$ is a feasible solution for the optimization problem solved by StageRoute (Eq. equation 3 applied to $\mathcal{M}_{\tau_{k}}$).
@@
Substituting these into Eq. equation 11:
@@
This bounds the true performance of the distribution $d^{*}$ chosen by the algorithm. Now we relate this to $V_{k}=V(b,\mathcal{D}_{k})$, the optimal performance within the chosen set $\mathcal{D}_{k}=\textnormal{supp}(d^{*})\subseteq\mathcal{M}_{\tau_{k}}$. The distribution $d^{*}$ is supported on $\mathcal{D}_{k}$ and uses at most $M_{\max}$ models (due to the constraint in Eq. equation 3). We examine its feasibility w.r.t. the true budget constraint. On $\mathcal{E}_{k}$:
@@
The term $d^{*}_{k,m}$ represents the selection weight for model $m$ in the deployment optimization at stage $k$. Thus, $\sum_{m\in\mathcal{D}_{k}}C_{m}(\tau_{k})d^{*}_{k,m}$ is a weighted average of the combined confidence radii for the deployed models. We are considering the case where the model has already been selected in the previous k1k-1 stages. Thus, we can bound the sum of $T_{k}(V^{*}_{k}-V_{k})$ by terms related to $\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}C_{m}(\tau_{k})$. Specifically,
@@
The StageRoute algorithm employs UCBs for rewards ($\mu^{U}_{m}$) and LCBs for costs ($c^{L}_{m}$) in its DeployOPT phase (Eq. equation 3). For a model $m$ that is newly available at stage $k$ (i.e., $t_{m}\leq\tau_{k}$ and its count of previous selections $N_{m}(\tau_{k})=0$), its initial empirical averages $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ are set based on priors. The confidence radius $f_{rad}(v,N_{m}(\tau_{k})+1)=\sqrt{\frac{\gamma v}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}$ becomes large for $N_{m}(\tau_{k})=0$. Specifically, with $N_{m}(\tau_{k})+1=1$:
@@
The total deployment regret bound in Lemma 8 highlights two distinct challenges in the deployment phase. The terms $\mathcal{O}(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})})$ and $\mathcal{O}(M_{\max}K\log(NT/\delta))$ capture the cost of learning the parameters of models that are considered for deployment. This cost depends on the horizon $T$, the number of deployment slots $M_{\max}$, the number of stages $K$, and logarithmic factors related to the total number of models $N$ and confidence $\delta$. The term $\mathcal{O}(NT/(M_{\max}K))$ reflects the structural cost imposed by the discovery bottleneck: when the universe of models $N$ is large compared to $M_{\max}$ and the number of adaptation opportunities $K$, there is an inherent regret incurred in sequentially exploring models to identify the best ones. This term can dominate if $N$ is very large or $K$ is small, underscoring the importance of the deployment frequency and capacity in dynamic LLM environments.
C.4 Bounding the Total Routing Regret
We now bound the total routing regret term $\mathcal{R}_{\textnormal{routing}}(T)$ identified in Lemma 4. This involves summing the per-stage regrets incurred by the Request Routing Phase of StageRoute due to using estimated model parameters within the deployed sets $\mathcal{D}_{k}$. This part of the proof follows a similar structure to the analysis of UCB-style algorithms for the Bandits with Knapsacks problem Agrawal and Devanur (2014).
@@ C.4 Bounding the Total Routing Regret
The proof follows the structure of Lemma 4.4 in Babaioff et al. (2015), adapted to our notation for models $m\in\mathcal{D}_{k}(\mathcal{A})$ and performance $\mu_{m}$, UCB $\mu^{U}_{m}(t)$, chosen model $m_{t}$, routing distribution $p_{t}^{*}$ (from Eq. equation 7), and summing over $t\in[\tau_{k},\tau_{k+1}-1]$ (length $T_{k}$).
@@
[…]
@@
The proof mirrors that of Lemma 9 (and Lemma 4.5 in Babaioff et al. (2015)), replacing performance with cost, $\mu_{m}$ with $\mathbb{E}[c_{m}]$, $\mu^{U}_{m}(t)$ with $c^{L}_{m}(t)$, and $r_{t}$ with $c_{t}$. The probability distribution $p_{t}^{*}$ is from Eq. equation 7. Key steps involve bounding $|\sum c_{t}-\sum\mathbb{E}[c_{m_{t}}]|$, $|\sum(\sum_{m}c^{L}_{m}(t)p_{t}^{*}(m))-\sum c^{L}_{m_{t}}(t)|$, and $|\sum\mathbb{E}[c_{m_{t}}]-\sum c^{L}_{m_{t}}(t)|$. The last term is bounded similarly using Lemma 3 and Lemma 2:
@@
From the algorithm’s choice of $p_{t}^{*}$ (solving Eq. equation 7) and property (1):
@@
Taking expectation conditioned on $\mathcal{D}_{k}$ (and implicitly on $\mathcal{E}$ for the bounds to hold), the result follows. The assumption $M_{k}\gamma\leq\mathcal{O}(B_{k})$ is used in concentration bounds for costs. Moreover, since the estimation errors for both performance and cost are governed by the same concentration inequalities, the expected budget violation is on the same order as the regret. ∎
@@
C.5 Total Regret Bound